# Echo chamber effects in signed networks

### v1

@@ Abstract
Echo chamber effects in social networks are generally attributed to the prevalence of interactions among like-minded peers. However, recent evidence has emphasized the role of hostile interactions between opposite-minded groups. Here, we model information propagation between such groups by generalizing popular contagion models to signed networks. We show that echo chambers spontaneously emerge in balanced networks, and in antibalanced ones for specific parameters. The robustness of our results is confirmed through simulations on various network topologies, including a real-world dataset.
@@
In recent years however, this evidence has been increasingly challenged [9, 10, 11, 12, 13, 14]. Scholars have pointed out that there is communication between opposite-minded communities, often in the form of hostile interactions [15, 16, 17]. Opposite-minded users may frame the same piece of information under different lights, reinforcing ideological divides while still allowing information to travel through partisan lines. Negative relations might be even more prevalent than positive ones for some users, such as journalists [16]. Thus, it is primordial that we distinguish between friends and foes when studying online debates [18].
@@
The presence of hostile interactions among individuals is often modelled as a signed network with positive and negative links [19]. The inclusion of negative links gives rise to novel phenomena, such as structural balance [20, 21]. A signed network is perfectly balanced if it can be divided into two groups, so that all links between agents of the same group are positive and all links between agents of different groups are negative [20]. In other words, it contains two groups that are antagonistic towards one another [22, 23]. While perfectly balanced networks are a mathematical idealization, studies have consistently shown a tendency towards structural balance in empirical networked systems of various types [24, 25, 26, 27]. On the other hand, when all links within each group are mostly negative and links between groups are mostly positive, the network is said to be antibalanced [28]. While not as common as balanced networks, antibalanced networks can emerge in historical, social and political contexts, like international relations during the Thirty Years’ War [29] or the U.S. Congress during the Carter and Reagan administrations (see [22] and the Appendix, Sec. D).
@@
These findings have motivated the introduction of the notion of degree of balance, with indices based on triangles [27, 22], random walks [30, 31, 32] or frustration metrics [33, 34], among others. Recent research has discovered intriguing connections between the degree of balance of social and political networks and their tendency towards polarization [22, 34]. The degree of balance also impacts the evolution of various dynamical processes [35, 36, 32], including contagion models [37, 38, 39], random walks [40], diffusive dynamics [41] and the voter model [42]. Conversely, structural balance can spontaneously arise as a result of opinion dynamics on signed networks [43, 44].
@@
Within this context, our contribution is two-fold. First, we extend to signed networks two popular models of information propagation, the Independent Cascade Model and the Linear Threshold Model [45, 46]. The key idea driving these extensions is that a single piece of information may be framed in different ways between, and within, antagonistic groups. Then, we use those models to analyze the key role played by the degree of balance in the emergence of echo chambers. We uncover a sudden transition as the degree of balance becomes positive, and an anomalous region where echo chambers can emerge in networks with negative degree of balance. Echoing empirical observations on the correlation between activity and polarisation [47, 48, 1, 49], we find reinforced echo chambers when information spreads more easily. The robustness of our results is confirmed on different network topologies, including a real-life dataset. As a by-product, we also provide a simple procedure to transform an unsigned network into a signed one with a prescribed degree of balance.
Signed social networks.
@@
We consider an undirected signed network with $N$ agents. Each agent is assigned a latent group membership, which represents the divergent opinions within the population. If all in-group interactions are friendly and all out-group interactions are hostile, then the network is perfectly balanced. In the opposite case, where in-groups are always hostile towards each other and friendly with out-groups, we call the network perfectly antibalanced [28]. To interpolate between these two extremes, we employ the triadic degree of balance $\tau$, commonly used in the literature [50, 27, 51, 22]. It is defined as:
@@
The majority of our simulations are run on networks with an underlying Erdös-Rényi (ER) topology with connection probability $p$. Hence, they do not exhibit communities from an unsigned point of view. Nevertheless, group memberships can induce mesoscopic structures affecting the degree of balance. We thus attribute signs to the links following a procedure that guarantees a pre-specified triadic degree of balance on average. This procedure is inspired by [53] and detailed in the Appendix, Sec. A.
Information propagation.
@@
We assume that an item is propagating throughout the network: a journal article, a tweet, a picture, etc. To describe its propagation, we adapt the archetypal Independent Cascade Model of simple contagion [45]. In this model, agents are initially all inactive (state 0), except for a few seeds who are the original emitters of the item (state 1). Their neighbors are thus exposed to the item, and may choose to share it themselves. Formally, in consecutive steps, each agent that was activated in the previous step activates each of its neighbors individually with probability $\lambda$. This probability is called the activation rate. If all attempted activations in a single step fail, the process stops. To account for signed links, we add a novel feature to these dynamics and propose the Signed Independent Cascade Model (SICM). Active agents can now be in one of two states $\{-1,+1\}$, reflecting two opposite-minded ways to frame the information: a fresh government decision may spark praise from some and criticism from others. When an agent is activated through a positive (resp. negative) link, they adopt the same (resp. opposite) state of the agent responsible for their activation. Once an agent is active, its state is fixed and never changes. More details can be found in the Appendix, Sec B.2.
@@
The SICM relies on a simple contagion mechanism: one exposure to information is enough to trigger a potential activation. In contrast, complex contagion mechanisms necessitate multiple exposures. This creates a different phenomenology [54, 55, 56] that has been argued to better model realistic behavioral dynamics [57, 58, 59]. For this purpose, we propose the Signed Linear Threshold Model (SLTM), adapted from [45, 46]. Similarly as for the SICM, agents can be in one of two states $\{-1,+1\}$ depending on who triggered their activation. Agents activate if the aggregated influence of their neighbors exceeds a threshold $T$—modelling details can be found in the Appendix, Sec. B2. While most of our results are presented for the SICM, we show at the end that the SLTM exhibits very similar behaviors.
Echo chamber effect
@@
In the Appendix, Sec. C.2. we prove that $E$ is a second-order moment; thus, we can interpret $E$ as a measure of the correlation between states and group memberships. The maximum $E=+1$ corresponds to perfect echo chambers, where everyone inside a group is in the same state, opposite to that of the other group. The minimum $E=-1$ is the case of an anti-echo chamber, where agents within the seed group disagree with the seeds, while agents of the opposite group agree with the seeds. When $E\approx 0$, there is no significant echo chamber effect. For the sake of conciseness, we refer to $E$ simply as the echo chamber effect, bearing in mind that it is actually a metric that quantifies this effect.
Results.
@@
Figure 1: Echo chamber effect on signed ER networks of various densities $p$. On the right half of the heatmaps are balanced networks $(\tau>0)$, on the left half are antibalanced networks $(\tau<0)$.
@@
Figure 2: Echo chamber effect in balanced ER networks. (a) $E$ as a function of the $\tau$ for various $\lambda$. (b) $E$ as a function of $\lambda$ for various $\tau$. (c) Activation rates $\lambda$ realising the minimum and maximum echo chamber effect for various $p$, averaged over all values of $\tau$. (d) Density $p$ realising the minimum and maximum echo chamber effect for various $\lambda$, averaged over all values of $\tau$. Shaded areas cover one standard deviation from the mean.
@@
We assess the echo chamber effect via simulations of the SICM on ER networks with $N=250$ agents. Figure 1 shows $E$ as function of $\tau$ and $\lambda$ on networks of various densities $p$. The difference between balanced ($\tau>0$) and antibalanced ($\tau<0$) networks is immediately striking, with a transition between the two that gets sharper as $p$ increases. In balanced networks, $E$ increases monotonically with $\tau$, as shown in Fig.2(a). Thus, we find that a stronger balance entails a higher echo chamber effect. In other words, echo chambers, much like polarization [22, 31] and dissensus [41], are amplified by structural balance in signed social networks.
@@
Furthermore, $E$ is in fact not monotonous with respect to $\lambda$, as shown in Fig. 2(b). There exists a nontrivial value of $\lambda$ for which $E$ is minimal, which decreases continuously with the network density (Fig. 2(c)). In contrast, the argmax of $E$ jumps discontinuously from $\lambda\approx 0$ to $\lambda\approx 1$. The same phenomenon is observed when computing the argmax of $E$ over $p$ (Fig.  2(d)). This suggests that minimizing the echo chamber effect is simpler than maximizing it: the optimal values of $\lambda$ and $p$ regarding minimization span a narrower range of values and are less susceptible to perturbations, whereas the optimal values for maximization vary widely between 00 and $1$ and can change abruptly in the critical region.
@@
In antibalanced networks, one could expect the same behavior, with $E$ decreasing monotonously towards $-1$ as $\tau$ becomes more negative. As we see in Fig. 1 however, this is not the case. Instead, we find that $E$ changes non-monotonously with both $p$ and $\lambda$ and $\tau$ (see Appendix, Sec. D). Specifically, $E$ reaches its minimum value, $E=-1$, both when $p$ and $\lambda$ are very high (e.g. $p=\lambda=1$) and when they are very low (e.g. $p=0.025,\lambda=0.01$). Perhaps most surprisingly, certain regions of the $(\tau<0,\lambda)$ half-plane display positive echo chamber effects. This anomalous region, corresponding to the red semi-circular area of Fig. 1, shifts downwards as the network density $p$ increases, and almost vanishes at high densities. Interestingly, Figure 3(a) reveals that for any given $\tau$, the boundaries of the anomalous region with respect to $p$ and $\lambda$ do not move. Instead, the value of $\tau$ only influences the strength of this anomaly.
@@
Figure 3: (a) Echo chamber effect as a function of network density and activation rate in antibalanced ER networks, for three different $\tau$. (b) Two realisations of the SICM on a signed ER network with $p=0.5,\tau=-1$, for $\lambda=0.08$ (echo chamber) and $\lambda=0.30$ (anti echo chamber). Circles correspond to agents of group $a$ and triangles to agents of group $b$. Agents are colored according to their state: grey for state 00, green for state $+1$ and pink for state $-1$.
@@
To better understand this phenomenon, we compare how the propagation unfolds for two different values of the activation rate. Figure 3(b) shows simulations of the SICM on a highly dense, perfectly antibalanced network ($\tau=-1$) with $\lambda=0.08$ and $\lambda=0.30$. Because $\tau=-1$, agents activated during odd steps adopt state $-1$ if they are in the seed group and state $+1$ otherwise, while the opposite happens during even steps. For $\lambda=0.30$, most activations occur in the first step, leading the seeding group to predominantly adopt state $-1$, resulting in a negative echo chamber effect. However, with $\lambda=0.08$, fewer activations happen initially, and most agents are activated in the second step. This causes the seeding group to adopt state $+1$, and generates a positive echo chamber effect even though the network is perfectly antibalanced. This alternance mechanism leads us to believe that this phenomenon is due to the synchronous aspect of the propagation process. Indeed, we show in the Appendix, Sec. D that these anomalies disappear with asynchronous activations. Additionally, the exact shape of the anomalous regions seem to change slightly with the network size (Appendix, Sec. D).
Robustness.
To demonstrate the robustness of our approach, we show that similar results are achieved (i) on networks with different topologies, and (ii) with our complex contagion model. In Fig. 4, we apply the SICM on a collection of networks with similar densities but various topological features. The Stochastic Block Model (SBM) generates networks with community structure, the model of Watts-Strogatz (WS) generates networks with clustering, and the model of Barabasi-Albert (BA) generates scale-free networks. Additionally, we consider a network extracted from a real-life dataset and exhibiting all of the above, Polblogs [61] (details in the Appendix, Sec. D). The echo chamber effect emerges when $\tau>0$ for every network, confirming that structural balance favors echo chambers regardless of the topology. Moreover, the anomalous region is also present for every network except the Watts-Strogatz one (Fig. 4f), due to clustering hindering the dominance of even-length paths.

### v2

# Modelling echo chamber effects in signed networks

@@ Abstract
Echo chamber effects in social networks are generally attributed to the prevalence of interactions among like-minded peers. However, recent evidence has emphasized the role of hostile interactions between opposite-minded groups. We investigate the role of polarization, identified with structural balance, in the formation of echo chambers in signed networks. To do so, we generalize the Independent Cascade Model and the Linear Threshold Model to describe information propagation in presence of negative edges. Antagonistic connections do not disrupt the flow of information, but instead, alter the way information is framed. Our results show that echo chambers spontaneously emerge in balanced networks, but also in antibalanced ones for specific parameters. This highlights that structural polarization and echo chambers do not necessarily display a one-to-one correspondence, showing instead a complex and often counterintuitive interplay. The robustness of our results is confirmed with a complex contagion model and through simulations in different network topologies, including real-world datasets.
@@ I Introduction
In recent years however, this evidence has been increasingly challenged [9, 10, 11, 12, 13, 14]. Scholars have pointed out that there is communication between opposite-minded communities, often in the form of hostile interactions [15, 16, 17]. Opposite-minded users may frame the same piece of information under different lights, reinforcing ideological divides while still allowing information to travel through partisan lines. Negative relations might be even more prevalent than positive ones for some users, such as journalists [16]. Similarly, recent research has highlighted the role of affective polarization, a phenomenon by which people experience positive feelings towards like-minded peers and negative feelings towards others [18, 19]. Thus, it is primordial that we distinguish between friends and foes when studying online debates [20].
@@
The presence of hostile interactions among individuals is often modelled as a signed network with positive and negative links [21]. The inclusion of negative links gives rise to novel phenomena, such as structural balance [22, 23]. A signed network is perfectly balanced if it can be divided into two groups, so that all links between agents of the same group are positive and all links between agents of different groups are negative [22]. In other words, it contains two groups that are antagonistic towards one another [24, 25]. While perfectly balanced networks are a mathematical idealization, studies have consistently shown a tendency towards structural balance in empirical networked systems of various types [26, 27, 28, 29]. On the other hand, when all links within each group are mostly negative and links between groups are mostly positive, the network is said to be antibalanced [30]. While not as common as balanced networks, antibalanced networks can emerge in historical, social and political contexts, like international relations during the Thirty Years’ War [31] or the U.S. Congress during the Carter and Reagan administrations (see [24] and Fig. 8(a)).
@@
These findings have motivated the introduction of the notion of degree of balance, with indices based on triangles [29, 24], random walks [32, 33, 34] or frustration metrics [35, 36], among others. Recent research has discovered intriguing connections between the degree of balance of social and political networks and their tendency towards affective [36] and ideological [24] polarization. In essence, the degree of balance—an inherently structural property—is strongly tied to the social and political phenomenon of polarization, a connection we term structural polarization. The degree of balance also impacts the evolution of various dynamical processes [37, 38, 34], including contagion models [39, 40, 41], random walks [42], diffusive dynamics [43] and the voter model [44]. Conversely, structural balance can spontaneously arise as a result of opinion dynamics on signed networks [45, 46].
@@
In this work, we add onto this literature by investigating the role of the degree of balance in the emergence of echo chambers. Doing so, we establish a correspondence between structural polarization and echo chamber effects, two distinct yet closely related phenomena. Our contribution is two-fold. First, we extend to signed networks two popular models of information propagation, the Independent Cascade Model and the Linear Threshold Model [47, 48]. The key idea driving these extensions is that a single piece of information may be framed in different ways between, and within, antagonistic groups. Second, we use those models to analyze the key role played by the degree of balance in the emergence of echo chambers. Echoing empirical observations on the correlation between activity and polarisation [49, 50, 1, 51], we find reinforced echo chambers in structurally polarized regimes where information spreads more easily. However, we also find echo chambers in regimes that do not correspond with polarized systems. This suggests that, contrary to what is commonly assumed, polarization and echo chambers are not equivalent concepts. The robustness of our results is confirmed on different network topologies, including two real-life datasets, a blog citation network [52] and a collection of U.S. congress bill co-sponsorship networks [24]. As a by-product, we also introduce a simple procedure to transform an unsigned network into a signed one with a prescribed degree of balance.
II Modeling
II.1 Signed social networks
@@
We consider an undirected signed network with $N$ agents. Each agent is assigned a latent group membership, which represents the divergent opinions within the population. If all in-group interactions are friendly and all out-group interactions are hostile, then the network is perfectly balanced. In the opposite case, where in-groups are always hostile towards each other and friendly with out-groups, we call the network perfectly antibalanced [30]. To interpolate between these two extremes, we employ the triadic degree of balance $\tau$, also called triadic influence [53]. This index has been commonly used in the literature to quantify the level of (structural) polarization in a given social system [54, 24, 29]. It is defined as:
@@
The majority of our simulations are run on networks with an underlying Erdös-Rényi (ER) topology with connection probability $p$. Hence, they do not exhibit communities from an unsigned point of view. Nevertheless, group memberships can induce mesoscopic structures affecting the degree of balance. We thus attribute signs to the links following a procedure that guarantees a pre-specified triadic degree of balance on average, inspired by [56]. To each link whose endpoints fall within the same group, we attribute a negative sign with probability $\eta$ and a positive sign otherwise; whereas links whose endpoints fall in different groups are positive with probability $\eta$ and negative otherwise. The parameter $\eta$ regulates the degree of balance in the network. For $\eta=0$ the network is (perfectly) balanced, for $\eta=1/2$ link signs are uncorrelated with group memberships, while for $\eta=1$ we obtain an antibalanced network. Intermediate values allow us to interpolate between the different degrees of balance.
To generate networks with a specific triadic degree of balance $\tau$, we make use of the following nonlinear relation, proven in the Supplementary Material [57, Sec. IIA]:
${\tau}=(1-2\eta)^{3}$ (2)
Figure 1: Degree of balance $\tau$ as a function of $\eta$, for various networks: ER with density $p=0.1$, ER with density $p=1.0$, BA with average degree $k=3$, and Polblogs [52] (details of the Polblogs dataset in Sec. III.2). Each point is averaged over 10 random assignments of link signs. Vertical bars indicate one standard deviation from the mean.
II.2 Information propagation
@@
We assume that an item is propagating throughout the network: a journal article, a tweet, a picture, etc. To describe its propagation, we adapt the archetypal Independent Cascade Model of simple contagion [47]. In this model, agents are initially all inactive (state 0), except for a few seeds who are the original emitters of the item (state 1). Their neighbors are thus exposed to the item, and may choose to share it themselves. Formally, in consecutive steps, each agent that was activated in the previous step activates each of its neighbors individually with probability $\lambda$. This probability is called the activation probability. If all attempted activations in a single step fail, the process stops. To account for signed links, we add a novel feature to these dynamics and propose the Signed Independent Cascade Model (SICM). Active agents can now be in one of two states $\{-1,+1\}$, reflecting two opposite-minded ways to frame the information: a fresh government decision may spark praise from some and criticism from others. When an agent is activated through a positive (resp. negative) link, they adopt the same (resp. opposite) state of the agent responsible for their activation. Once an agent is active, its state is fixed and never changes. Technical details of the implementation can be found in the Supplementary Material [57, Sec. IA].
@@
The SICM relies on a simple contagion mechanism: one exposure to information is enough to trigger a potential activation. In contrast, complex contagion mechanisms necessitate multiple exposures. This creates a different phenomenology [58, 59, 60] that has been argued to better model realistic behavioral dynamics [61, 62, 63]. For this purpose, we propose the Signed Linear Threshold Model (SLTM), adapted from [47, 48]. Similarly as for the SICM, agents can be in one of two states $\{-1,+1\}$ depending on who triggered their activation. Agents activate if the aggregated influence of their neighbors exceeds a threshold $T$. Details of the modeling and implementation can be found in the Supplementary Material [57, Sec. IB]. While most of our results are presented for the SICM, we show in Sec. III.2 that the SLTM exhibits similar behaviors.
II.3 Echo chamber effect
@@
In the Supplementary Material [57, Sec. IIB] we prove that $E$ is a second-order moment; thus, we can interpret $E$ as a measure of the correlation between states and group memberships. The maximum $E=+1$ corresponds to perfect echo chambers, where everyone inside a group is in the same state, opposite to that of the other group. The minimum $E=-1$ is the case of an anti-echo chamber, where agents within the seed group disagree with the seeds, while agents of the opposite group agree with the seeds. When $E\approx 0$, there is no significant echo chamber effect. In other words, $E$ measures the level of homogeneity in the framing of information within and between opposite-minded groups, rather than the presence or absence of cross-cutting exchanges. For the sake of conciseness, we refer to $E$ simply as the echo chamber effect, bearing in mind that it is actually a metric that quantifies this effect.
III Results
III.1 Simple contagion on Erdös-Rényi networks
@@
Figure 2: Echo chamber effect $E$ on signed Erdös-Rényi (ER) networks of various densities $p$, for SICM dynamics and as a function of the activation probability $\lambda$ and the triadic balance $\tau$. On the right half of the heatmaps are balanced networks $(\tau>0)$, on the left half are antibalanced networks $(\tau<0)$. In this figure and the remaining ones, we set $N=250$ nodes and $M=800$ realizations per data point.
@@
Figure 3: Echo chamber effect $E$ in balanced ER networks for SICM dynamics. (a) $E$ as a function of the triadic balance $\tau$ for various activation probabilities $\lambda$. The inset on the right plot shows the match between simulation results (red circles) and theoretical prediction ($E=\tau^{1/3}$, blue line) for $\lambda=p=1$. (b) $E$ as a function of $\lambda$ for various $\tau$. (c) activation probability realising the minimum and maximum echo chamber effect for various $p$, averaged over all values of $\tau$. (d) Density $p$ realising the minimum and maximum echo chamber effect for various $\lambda$, averaged over all values of $\tau$. Shaded areas cover one standard deviation from the mean.
@@
We assess the echo chamber effect via simulations of the SICM on ER networks with $N=250$ agents. Figure 2 shows $E$ as function of $\tau$ and $\lambda$ on networks of various densities $p$. The difference between balanced ($\tau>0$) and antibalanced ($\tau<0$) networks is immediately striking, with a transition between the two that gets sharper as $p$ increases. In balanced networks, $E$ increases monotonically with $\tau$, as shown in Fig.3(a). Thus, we find that a stronger balance entails a higher echo chamber effect. In other words, echo chambers, much like polarization [24, 33] and dissensus [43], are amplified by the level of structural balance in balanced signed social networks. Note, however, that even in this simple scenario, the relation between the level in balance and the echo chamber effect is non-linear. In fact, in the Supplementary Material [57, Sec. IIC], we prove that for $\lambda=p=1$, this relation is given by $E=\tau^{1/3}$. See the inset in Fig.3(a) for illustration.
@@
In Fig. 3(b), we notice that $E$ is also not monotonous with respect to $\lambda$. There exists a nontrivial value of $\lambda$ for which $E$ is minimal, which decreases continuously with the network density (Fig. 3(c)). In contrast, the argmax of $E$ jumps discontinuously from $\lambda\approx 0$ to $\lambda\approx 1$. The same phenomenon is observed when computing the argmax of $E$ over $p$ (Fig.  3(d)). This suggests that minimizing the echo chamber effect is simpler than maximizing it: the optimal values of $\lambda$ and $p$ regarding minimization span a narrower range of values and are less susceptible to perturbations, whereas the optimal values for maximization vary widely between 00 and $1$ and can change abruptly in the critical region.
@@
In antibalanced networks, one could expect the same behavior, with $E$ decreasing monotonously towards $-1$ as $\tau$ becomes more negative. As we see in Fig. 2 however, this is not the case. Instead, we find that $E$ changes non-monotonously with both $p$ and $\lambda$ and $\tau$ (see the Supplementary Material [57, Sec. III]). Specifically, $E$ reaches its minimum value, $E=-1$, both when $p$ and $\lambda$ are very high (e.g. $p=\lambda=1$) and when they are very low (e.g. $p=0.025,\lambda=0.01$). This phenomenon explicitly challenges the identification of structural polarization with echo chambers. Perhaps most surprisingly, certain regions of the $(\tau<0,\lambda)$ half-plane display positive echo chamber effects. This anomalous region, corresponding to the red semi-circular area of Fig. 2, shifts downwards as the network density $p$ increases, and almost vanishes at high densities. Interestingly, Figure 4(a) reveals that for any given $\tau$, the boundaries of the anomalous region with respect to $p$ and $\lambda$ do not move. Instead, the value of $\tau$ only influences the strength of this anomaly.
@@
Figure 4: (a) Echo chamber effect as a function of network density and activation probability in antibalanced ER networks, for SICM dynamics and three different $\tau$. (b) Two realisations of the SICM on a signed ER network with $p=0.5,\tau=-1$, for $\lambda=0.08$ (echo chamber) and $\lambda=0.30$ (anti echo chamber). Circles correspond to agents of group $a$ and triangles to agents of group $b$. Agents are colored according to their state: grey for state 00, green for state $+1$ and pink for state $-1$.
@@
To better understand this phenomenon, we compare how the propagation unfolds for two different values of the activation probability. Figure 4(b) shows simulations of the SICM on a highly dense, perfectly antibalanced network ($\tau=-1$) with $\lambda=0.08$ and $\lambda=0.30$. Because $\tau=-1$, agents activated during odd steps adopt state $-1$ if they are in the seed group and state $+1$ otherwise, while the opposite happens during even steps. For $\lambda=0.30$, most activations occur in the first step, leading the seeding group to predominantly adopt state $-1$, resulting in a negative echo chamber effect. However, with $\lambda=0.08$, fewer activations happen initially, and most agents are activated in the second step. This causes the seeding group to adopt state $+1$, and generates a positive echo chamber effect even though the network is perfectly antibalanced. This alternance mechanism leads us to believe that this phenomenon is due to the synchronous aspect of the propagation process. Indeed, we show in the Supplementary Material [57, Sec. III] that these anomalies disappear with asynchronous activations. Additionally, the exact shape of the anomalous regions seem to change slightly with the network size (Supplementary Material [57, Sec. III]).
III.2 Robustness
To demonstrate the robustness of our approach, we show that similar results are achieved (i) on networks with different topologies, and (ii) with our complex contagion model. In Fig. 5, we apply the SICM on a collection of networks with similar densities but various topological features. The Stochastic Block Model (SBM) generates networks with community structure, the model of Watts-Strogatz (WS) generates networks with clustering, and the model of Barabasi-Albert (BA) generates scale-free networks. Additionally, we consider a network extracted from the Poblogs dataset [52], which exhibits all of the above. The dataset comprises an unsigned network of political blogs collected during the campaign of the 2004 U.S. presidential elections. Blogs split into two groups, liberal and conservative, and there is an edge between two blogs if one referred to the other at least once. We restrict ourselves to the largest connected component of the network, with 1,222 nodes and 19,021 links. Link signs are generated following the procedure described in Sec. II.1, to obtain a collection of networks with varying degrees of balance.